# Phase diagram of compressible and paired states in the quarter-filled Landau level

### v1

@@ Abstract
Quantum Hall plateaus at quarter fillings occur in GaAs wide quantum wells, hole-doped GaAs, and bilayer graphene. However, the interactions favoring incompressible states over compressible composite-Fermi liquids at such fillings are not well understood. We devise a method of computing the trial energies for Haldane pseudopotentials via Monte Carlo sampling. Applying it to the quarter-filled lowest Landau level, we find that tuning the third and fifth pseudopotential can stabilize anti-Pfaffian, Moore-Read, and f-wave states. The smallest deviations from pure Coulomb interactions are required by anti-Pfaffian, whose presence is indicated by daughter states in recent experiments of bilayer graphene at $\nu=\frac{3}{4}$.
@@ I Introduction and main result
The fractional quantum Hall effect Tsui et al. [1982], Laughlin [1983], Haldane [1983], Rammal et al. [1983] at even-denominators has been attracting widespread attention due to the possibility of hosting non-Abelian anyons. Wen [1991], Nayak et al. [2008], Moore and Read [1991], Read and Green [2000] The best-known candidate for such a state is the half-filled plateau in the first excited Landau level of GaAs, i.e., at the filling factor $\nu=\frac{5}{2}$illett et al. [1987] The gap at this filling factor is attributed to the pairing of composite fermions Read and Green [2000] – electrons bound to two flux quanta, which exactly cancel the electron charge at half-filling. Jain [1989, 2007] Numerical studies of the half-filled Landau level found that Coulomb interactions favor a compressible composite-Fermi liquid (CFL) in the lowest Landau level (0LL). In the first excited Landau level (1LL), a paired state is stabilized instead. Morf [1998], Rezayi and Haldane [2000] The transition from metallic to paired state can be induced by reducing the first Haldane pseudopotential, Haldane [1983] which suppresses the repulsion between electrons at short distances.
Further numerical studies Wójs et al. [2010], Storni et al. [2010], Feiguin et al. [2008, 2009], Peterson et al. [2008], Rezayi and Simon [2011], Pakrouski et al. [2015] confirmed these results and focused on distinguishing the leading candidates: the Moore-Read Moore and Read [1991] state and its particle-hole conjugate, known as anti-Pfaffian. Levin et al. [2007], Lee et al. [2007] However, neither of these states is compatible with thermal transport measurements in GaAs quantum wells at $\nu=\frac{5}{2}$anerjee et al. [2018], Dutta et al. [2022a, b], Melcer et al. [2022], Paul et al. [2023], Melcer et al. [2024] These experiments instead indicate a particle-hole symmetric topological order. Son [2015] Such a phase has never been found numerically, and disorder may be essential for understanding its origin. Mross et al. [2018], Wang et al. [2018], Lian and Wang [2018] Still, the finding that 1LL is conducive to pairing at half filling appears robust and extends to other quantum Hall platforms, such as bilayer graphene. Ki et al. [2014], Kim et al. [2015], Li et al. [2017], Zibrov et al. [2017], Assouline et al. [2024]
@@
Closely related paired states can occur in quarter-filled Landau levels. Read and Green [2000] Here, binding four flux quanta to each electron yields charge-neutral composite fermions. Plateaus at quarter filling have been observed in GaAs-wide quantum wells Luhman et al. [2008], Shabani et al. [2009] and, more recently, in hole-doped GaAs Wang et al. [2022, 2023] and bilayer graphene. Kumar et al. [2024] Surprisingly, all of those states arise in the 0LL, although Landau-level mixing may be significant. The conditions that favor pairing at quarter filling are less well understood than those at half filling. Exact diagonalization studies are limited to $N_{e}\leq 12$ particles, and the fate of these systems in the thermodynamic limit remains unclear. Papić et al. [2009a], Sharma et al. [2024] Still, Monte Carlo studies of large systems indicate that an f-wave paired stateWen [1991] may be favored in wide quantum wells at quarter filling. Papić et al. [2009a], Faugno et al. [2019], Sharma et al. [2024]
@@
In this work, we systematically analyze the competition between compressible and paired states in a quarter-filled Landau level with general interactions. To achieve this, we develop an efficient method for evaluating the trial wavefunction energies for arbitrary choices of Haldane pseudopotentials directly with Monte Carlo simulations. Our numerical findings are summarized by the phase diagram in Fig. 1. It is obtained by computing the energy of different trial states at system sizes up to $N_{e}=56$ particle and extrapolating to the thermodynamic limit. We find that the third and fifth Haldane pseudopotentials yield the strongest distinction between competing states at this filling. In the lowest Landau level with pure Coulomb interactions we find that the composite-Fermi liquid is favored. The paired anti-Pfaffian, Moore-Read, and f-wave states can all be stabilized by small changes in pseudopotentials; see Fig. 2. In particular, the anti-Pfaffian paired state is close to the Coulomb point in the 0LL. This finding is consistent with recent experimental results Kumar et al. [2024] of $\nu=\frac{3}{4}$ plateaus in bilayer graphene, where plateaus are accompanied by ‘daughter’ states Levin and Halperin [2009], Yutushui et al. [2024], Zheltonozhskii et al. [2024], Zhang et al. [2024] that support the anti-Pfaffian pairing.
@@
Figure 2: The overlaps of the ground state $|GS(x)\rangle$ of Eq. (26) with the trial states at $\nu=\frac{1}{4}$ for $N_{e}=10$. The angles $\theta_{\Psi}$ encode the directions in the $\delta H_{3}-\delta H_{5}$ plane indicated by arrows in Fig. 1, that favor the paired state $\Psi$. For the anti-Pfaffian and f-wave cases, modifying the third and fifth Haldane pseudopotentials along these directions results in transitions from compressible CFL state for pure 0LL Coulomb interactions $(x=0)$ to the paired state at positive $x$. At the Moore-Read shift, the ground state at $x=0$ has non-zero angular momentum. Since $N_{e}=10$ does not correspond to a filled shell Rezayi and Read [1994] of composite fermions at this shift, this finding is consistent with a composite Fermi liquid ground state. Here, we show the overlap of the trial state with the lowest zero angular momentum state $|GS_{L=0}(x)\rangle$, in addition to the one with the overall ground state. For all three shifts, the squared overlap with paired trial states peaks at positive $x$ and is above $90\%$.
@@
The rest of this work is organized as follows: In Sec. II, we briefly introduce the key players: (i) Haldane pseudopotentials, (ii) Real-space harmonics, and (iii) Static structure factors. Sec. III describes the numerical methods developed in this work. After providing a general mapping between Haldane pseudopotential and real-space harmonics, we introduce the refinements necessary to obtain numerically stable results. In Sec. IV.1, we benchmark this approach by comparing trial energies from Monte Carlo to those obtained via exact diagonalization. In Sec. IV.2, we present results for CFLs at different shifts, filling factors, and Landau levels. Sec. V contains our main results on paired states at half and quarter filling. We conclude in Sec. VI by discussing the implications and possible extensions of our work. Appendices contain additional details: In Appendix A, we derive the real-space form of higher Landau-level pseudopotentials. Appendix B discusses finite size correction and introduces the normalization of energies throughout the paper. Appendix C contains details of the Landau level projection. Appendix D lists overlaps of paired states and CFLs.
@@ II Energies in Exact Diagonalization and Monte Carlo
Arbitrary two-body interactions between particles that occupy a single Landau level can be efficiently parametrized by a set of Haldane pseudopotentials Haldane [1983] as
@@
In numerical studies, one frequently chooses a specific interaction potential, encoded via $\{H^{0}_{L}\}$, as a reference point and varies a small number of pseudopotentials $\{\delta H_{L}\}$ to obtain a phase diagram like the one in Fig. 1. With exact diagonalization, one can compute the expectation values of any Hamiltonian $\mathcal{H}$ in a state $\Psi$ as $E_{\Psi}=\sum_{L}H_{L}\langle\hat{\cal P}_{L}\rangle_{\Psi}$ójs and Quinn [2005] However, this technique is limited to small systems, and it is often challenging to extrapolate to the thermodynamic limit.
@@ II.1 Energies from Monte Carlo
Given a trial wavefunction $\Psi(\bm{r})$ and real-space potential $V(r)$, one can efficiently use Monte Carlo sampling to evaluate the integral
$E_{V}=\frac{1}{2}\int|\Psi(\{\bm{r}_{k}\})|^{2}\sum_{i\neq j}V(|{\bm{r}}_{i}-{\bm{r}}_{j}|)~{}.$ (2)
@@
The energy of higher Landau levels can be similarly computed if $\Psi$ is projected to the corresponding Landau level. However, most trial states and known methods of Landau-level projection are restricted to the lowest Landau level. The different single-particle orbitals in the $n$th Landau level compared to 0LL can alternatively be incorporated by a modified interaction. Approximate real-space potentials $V^{\text{$n$LL}}_{\text{approx}}(r)$ for this purpose have been constructed, e.g., in Refs. Park et al., 1998, Töke et al., 2005, by requiring the lowest pseudopotentials to coincide. The price to pay is that $V^{\text{$n$LL}}_{\text{approx}}(r)$ is less smooth, and the Monte Carlo algorithm takes a longer time to converge. Thus, only the first few Landau levels are accessible.
@@ II.2 Static structure factors and energy
$G(\theta)=\frac{2}{N^{2}_{e}}\int|\Psi(\{\bm{r}_{i}\})|^{2}\sum_{i\neq j}\delta({\bm{r}}_{i}\cdot{\bm{r}}_{j}-\cos\theta),$ (4)
@@
$G_{k}=\frac{2k+1}{N_{e}^{2}}\sum_{i\neq j}\left\langle P_{k}({\bm{r}}_{i}\cdot{\bm{r}}_{j})\right\rangle.$ (8)
@@
which can diagnose metallic states via singular behavior around $k=2k_{F}\approx\sqrt{N_{e}}$amilla et al. [1997], Wójs and Quinn [2005]
Figure 3: (a) The harmonics $V_{k}/(2k+1)$ encoding Coulomb interactions in the 1LL grow as $k^{3}$. The refined harmonics $\tilde{V}_{k}$ of Sec. III.1 instead decay to zero. Different choices of the refinement parameter $\alpha$ in Eq. (19) yield very similar coefficients. This behavior is shown here for $l=87$ but is representative of general system sizes. (b) Expansion coefficients $G_{k}$ of the density-density correlation function obtained via Monte Carlo for a Laughlin state with $N_{e}=30$. The coefficients are close to zero for $k\gtrsim 15$ but exhibit noise due to finite sampling, as seen in the inset. Without refinement, this noise is significantly amplified by the growing coefficient $V_{k}$ in Eq. (7). After refinement, the harmonics $\tilde{V}_{k}$ approach zero for large $k$, resulting in a numerically stable estimate for the energy.
@@
where the $12q+$ 1-dimensional transition matrix Wooten [2013] is
@@ III.1 Refined interactions
Above, we identified the growth of $V_{k}$ at large $k$ as the source of numerical instability. Our resolution to this problem is to tune the even $H_{L}$ to suppress $V_{k}$ with large $k$. Given a real-space potential expressed through $V_{k}$, we find the corresponding Haldane pseudopotentials $H_{L}$. Then, we divide $H_{L}$ into odd ${\bm{H}}_{\text{odd}}=(H_{1},H_{3},\ldots)$ and even ${\bm{H}}_{\text{even}}=(H_{0},H_{2},\ldots)$ parts written as vectors of dimensions $d_{\text{odd}}=\lfloor\frac{2q+1}{2}\rfloor$ and $d_{\text{even}}=\lfloor q\rfloor+1$, respectively. We now replace ${\bm{H}}_{\text{even}}$ with a vector $\tilde{{\bm{H}}}_{\text{even}}$ of free parameters and introduce the equivalent harmonics as
$\tilde{{\bm{V}}}(\tilde{{\bm{H}}}_{\text{even}})=N(q)\cdot\tilde{{\bm{H}}},$ (18)
with the matrix $N(q)$ specified in Eq. (15). The vector $\tilde{{\bm{H}}}=\tilde{{\bm{H}}}_{\text{even}}\cup{\bm{H}}_{\text{odd}}$ is comprised of the original odd pseudopotentials ${\bm{H}}_{\text{odd}}$ and the free parameters $\tilde{{\bm{H}}}_{\text{even}}$. By construction, $\tilde{{\bm{V}}}$ yields the same energy as ${{\bm{V}}}$.
To reduce statistical noise, we tune $\tilde{{\bm{H}}}_{\text{even}}$ so that $\tilde{V}_{k}$ decay quickly by minimizing the cost function
${\cal F}(\tilde{{\bm{H}}}_{\text{even}})=\sum_{k=1}^{2q}[\tilde{V}_{k}(\tilde{{\bm{H}}}_{\text{even}})]^{2}(2k+1)^{\alpha},$ (19)
with $\alpha\geq 2$ to penalize large $k$ harmonics over small $k$foo We take $\alpha=4$ for the data presented in this work, but we find the final result to be largely insensitive of $\alpha$; see Fig. 3. After determining $\tilde{{\bm{H}}}^{\text{opt}}_{\text{even}}$ that minimizes ${\cal F}(\tilde{{\bm{H}}}_{\text{even}})$, we compute the refined values of the harmonics $\tilde{V}_{k}(\tilde{{\bm{H}}}^{\text{opt}}_{\text{even}})$. Using them to compute the energy Eq. (7) yields the same value as using the original $V_{k}=\tilde{V}_{k}({{\bm{H}}}_{\text{even}})$ but with less noise. The optimization procedure can be done semi-analytically, e.g., with Mathematica, and is not numerically expensive for any system size.
We note that Eq. (7) assumes spherically symmetric states. Nevertheless, the refined harmonics $\tilde{{\bm{V}}}$ can be cast into a real-space form $\tilde{V}(\theta)$ with Eq. (6) that can be used to compute the energy of non-zero angular momentum states within Monte Carlo simulations.

### v3

@@ Abstract
Quantum Hall plateaus at quarter fillings occur in GaAs wide quantum wells, hole-doped GaAs, and bilayer graphene. However, the interactions favoring incompressible states over compressible composite-Fermi liquids at such fillings are not well understood. We devise a method for the computation of the trial energies for Haldane pseudopotentials via Monte Carlo sampling. Applying it to the quarter-filled lowest Landau level, we find that tuning the third and fifth pseudopotential can stabilize anti-Pfaffian, Moore-Read, and f-wave states. The smallest deviations from pure Coulomb interactions are required by anti-Pfaffian, whose presence is indicated by daughter states in recent experiments of bilayer graphene at $\nu=\frac{3}{4}$.
@@ I Introduction and main result
The fractional quantum Hall effect Tsui et al. (1982); Laughlin (1983); Haldane (1983); Rammal et al. (1983) at even-denominators has been attracting widespread attention due to the possibility of hosting non-Abelian anyons. Wen (1991); Nayak et al. (2008); Moore and Read (1991); Read and Green (2000) The best-known candidate for such a state is the half-filled plateau in the first excited Landau level of GaAs, i.e., at the filling factor $\nu=\frac{5}{2}$illett et al. (1987) The gap at this filling factor is attributed to the pairing of composite fermions Read and Green (2000) – electrons bound to two flux quanta, which exactly cancel the electron charge at half-filling. Jain (1989a, 2007) Numerical studies of the half-filled Landau level found that Coulomb interactions favor a compressible composite-Fermi liquid (CFL) in the lowest Landau level (0LL). In the first excited Landau level (1LL), a paired state is stabilized instead. Morf (1998); Rezayi and Haldane (2000) The transition from metallic to paired state can be induced by reducing the first Haldane pseudopotential, Haldane (1983) which suppresses the repulsion between electrons at short distances.
Further numerical studies Wójs et al. (2010); Storni et al. (2010); Feiguin et al. (2008, 2009); Peterson et al. (2008); Rezayi and Simon (2011); Pakrouski et al. (2015) confirmed these results and focused on distinguishing the leading candidates: the Moore-Read Moore and Read (1991) state and its particle-hole conjugate, known as anti-Pfaffian. Levin et al. (2007); Lee et al. (2007) However, neither of these states is compatible with thermal transport measurements in GaAs quantum wells at $\nu=\frac{5}{2}$anerjee et al. (2018); Dutta et al. (2022a, b); Melcer et al. (2022); Paul et al. (2024); Melcer et al. (2024); Paul et al. (2024); Dutta et al. (2022c) These experiments instead indicate a particle-hole symmetric topological order. Son (2015) Such a phase has never been found numerically, and disorder may be essential for understanding its origin. Mross et al. (2018); Wang et al. (2018); Lian and Wang (2018) Still, the finding that 1LL is conducive to pairing at half filling appears robust and extends to other quantum Hall platforms, such as bilayer graphene. Ki et al. (2014); Kim et al. (2015); Li et al. (2017); Zibrov et al. (2017); Assouline et al. (2024)
@@
Closely related paired states can occur in quarter-filled Landau levels. Read and Green (2000) Here, binding four flux quanta to each electron yields charge-neutral composite fermions. Plateaus at quarter filling have been observed in GaAs-wide quantum wells Luhman et al. (2008); Shabani et al. (2009) and, more recently, in hole-doped GaAs Wang et al. (2022, 2023) and bilayer graphene. Kumar et al. (2024) Surprisingly, all of those states arise in the 0LL, although Landau-level mixing may be significant. The conditions that favor pairing at quarter filling are less well understood than those at half filling. Exact diagonalization studies are limited to $N_{e}\leq 12$ particles, and the fate of these systems in the thermodynamic limit remains unclear. Papić et al. (2009a); Sharma et al. (2024) Still, Monte Carlo studies of large systems indicate that an f-wave paired stateWen (1991) may be favored in wide quantum wells at quarter filling. Papić et al. (2009a); Faugno et al. (2019); Sharma et al. (2024) Fixed-phase diffusion Monte-Carlo approach finds that anti-Pfaffian can be stabilized in the presence of strong Landau level mixing. Zhao et al. (2023)
@@
In this work, we systematically analyze the competition between compressible and paired states in a quarter-filled Landau level with general interactions. To achieve this, we develop an efficient method for evaluating the trial wavefunction energies for arbitrary choices of Haldane pseudopotentials directly with Monte Carlo simulations. Our numerical findings are summarized by the phase diagram in Fig. 1. It is obtained by computing the energy of different trial states at system sizes up to $N_{e}=56$ particle and extrapolating to the thermodynamic limit. We find that the third and fifth Haldane pseudopotentials yield the strongest distinction between competing states at this filling. In the lowest Landau level with pure Coulomb interactions we find that the composite-Fermi liquid is favored. The paired anti-Pfaffian, Moore-Read, and f-wave states can all be stabilized by small changes in pseudopotentials; see Fig. 2. In particular, the anti-Pfaffian paired state is close to the Coulomb point in the 0LL. This finding is consistent with recent experimental results Kumar et al. (2024) of $\nu=\frac{3}{4}$ plateaus in bilayer graphene, where plateaus are accompanied by ‘daughter’ states Levin and Halperin (2009); Yutushui et al. (2024a); Zheltonozhskii et al. (2024); Zhang et al. (2024) that support the anti-Pfaffian pairing.
@@
Figure 2: The overlaps of the ground state $|GS(x)\rangle$ of Eq. (26) with the trial states at $\nu=\frac{1}{4}$ for $N_{e}=10$. The angles $\theta_{\Psi}$ encode the directions in the $\delta H_{3}-\delta H_{5}$ plane indicated by arrows in Fig. 1, that favor the paired state $\Psi$. For the anti-Pfaffian and f-wave cases, modifying the third and fifth Haldane pseudopotentials along these directions results in transitions from compressible CFL state for pure 0LL Coulomb interactions $(x=0)$ to the paired state at positive $x$. At the Moore-Read shift, the ground state at $x=0$ has non-zero angular momentum. Since $N_{e}=10$ does not correspond to a filled shell Rezayi and Read (2005) of composite fermions at this shift, this finding is consistent with a composite Fermi liquid ground state. Here, we show the overlap of the trial state with the lowest zero angular momentum state $|GS_{L=0}(x)\rangle$, in addition to the one with the overall ground state. For all three shifts, the squared overlap with paired trial states peaks at positive $x$ and is above $90\%$.
@@
The rest of this work is organized as follows: In Sec. II, we briefly introduce the key players: (i) Haldane pseudopotentials, (ii) Real-space harmonics, and (iii) Static structure factors. Sec. III describes the numerical methods developed in this work. After providing a general mapping between Haldane pseudopotential and real-space harmonics, we introduce the refinements necessary to obtain numerically stable results. In Sec. IV.1, we benchmark this approach by comparing trial energies from Monte Carlo to those obtained via exact diagonalization. In Sec. IV.2, we present results for CFLs at different shifts, filling factors, and Landau levels. Sec. V contains our main results on paired states at half and quarter filling. We conclude in Sec. DATA AVAILABILITY by discussing the implications and possible extensions of our work. Appendices contain additional details: In Appendix A, we derive the real-space form of higher Landau-level pseudopotentials. Appendix B discusses finite size correction and introduces the normalization of energies throughout the paper. Appendix D contains details of the Landau level projection. Appendix E lists overlaps of paired states and CFLs.
@@ II Energies in Exact Diagonalization and Monte Carlo
Arbitrary two-body interactions between particles that occupy a single Landau level can be efficiently parametrized by a set of Haldane pseudopotentials Haldane (1983) as
@@
In numerical studies, one frequently chooses a specific interaction potential, encoded via $\{H^{0}_{L}\}$, as a reference point and varies a small number of pseudopotentials $\{\delta H_{L}\}$ to obtain a phase diagram like the one in Fig. 1. With exact diagonalization, one can compute the expectation values of any Hamiltonian $\mathcal{H}$ in a state $\Psi$ as $E_{\Psi}=\sum_{L}H_{L}\langle\hat{\cal P}_{L}\rangle_{\Psi}$ Wójs and Quinn (2005); see also Ref. Jain, 2007, in particular Appendix F for a review of the exact diagonalization method. However, this technique is limited to small systems, and it is often challenging to extrapolate to the thermodynamic limit.
@@ II.1 Energies from Monte Carlo
Given a trial wavefunction $\Psi(\bm{r})$ and real-space potential $V(r)$, one can efficiently use Monte Carlo sampling (see Appending K of Ref. Jain, 2007 for details) to evaluate the integral
$E_{V}=\frac{1}{2}\int|\Psi(\{\bm{r}_{k}\})|^{2}\sum_{i\neq j}V(|\vec{r}_{i}-\vec{r}_{j}|)~{}.$ (2)
@@
The energy of higher Landau levels can be similarly computed if $\Psi$ is projected to the corresponding Landau level. However, most trial states and known methods of Landau-level projection are restricted to the lowest Landau level. The different single-particle orbitals in the $n$th Landau level compared to 0LL can alternatively be incorporated by a modified interaction. Approximate real-space potentials $V^{\text{$n$LL}}_{\text{approx}}(r)$ for this purpose have been constructed, e.g., in Refs. Park et al., 1998; Töke et al., 2005, by requiring the lowest pseudopotentials to coincide. The price to pay is that $V^{\text{$n$LL}}_{\text{approx}}(r)$ is less smooth, and the Monte Carlo algorithm takes a longer time to converge. Thus, only the first few Landau levels are accessible.
@@ II.2 Static structure factors and energy
$G(\theta)=\frac{2}{N^{2}_{e}}\int|\Psi(\{\bm{r}_{i}\})|^{2}\sum_{i\neq j}\delta(\vec{r}_{i}\cdot\vec{r}_{j}-\cos\theta),$ (4)
@@
$G_{k}=\frac{2k+1}{N_{e}^{2}}\sum_{i\neq j}\left\langle P_{k}(\vec{r}_{i}\cdot\vec{r}_{j})\right\rangle.$ (8)
@@
which can diagnose metallic states via singular behavior around $k=2k_{F}\approx\sqrt{N_{e}}$amilla et al. (1997); Wójs and Quinn (2005)
Figure 3: (a) The harmonics $V_{k}/(2k+1)$ encoding Coulomb interactions in the 1LL grow as $k^{3}$. The refined harmonics $\tilde{V}_{k}$ of Sec. III.1 instead decay to zero. Different choices of the refinement parameter $\alpha$ in Eq. (19) yield very similar coefficients. This behavior is shown here for $872l=$ 87 but is representative of general system sizes. (b) Expansion coefficients $G_{k}$ of the density-density correlation function obtained via Monte Carlo for a Laughlin state with $N_{e}=30$. The coefficients are close to zero for $k\gtrsim 15$ but exhibit noise due to finite sampling, as seen in the inset. Without refinement, this noise is significantly amplified by the growing coefficient $V_{k}$ in Eq. (7). After refinement, the harmonics $\tilde{V}_{k}$ approach zero for large $k$, resulting in a numerically stable estimate for the energy.
@@
where the $12q+$ 1-dimensional transition matrix Wooten (2013) is
@@ III.1 Refined interactions
Above, we identified the growth of $V_{k}$ at large $k$ as the source of numerical instability. Our resolution to this problem is to tune the even $H_{L}$ to suppress $V_{k}$ with large $k$. Given a real-space potential expressed through $V_{k}$, we find the corresponding Haldane pseudopotentials $H_{L}$. Then, we divide $H_{L}$ into odd $\vec{H}_{\text{odd}}=(H_{1},H_{3},\ldots)$ and even $\vec{H}_{\text{even}}=(H_{0},H_{2},\ldots)$ parts written as vectors of dimensions $d_{\text{odd}}=\lfloor\frac{2q+1}{2}\rfloor$ and $d_{\text{even}}=\lfloor q\rfloor+1$, respectively. We now replace $\vec{H}_{\text{even}}$ with a vector $\tilde{\vec{H}}_{\text{even}}$ of free parameters and introduce the equivalent harmonics as
$\tilde{\vec{V}}(\tilde{\vec{H}}_{\text{even}})=N(q)\cdot\tilde{\vec{H}},$ (18)
with the matrix $N(q)$ specified in Eq. (15). The vector $\tilde{\vec{H}}=\tilde{\vec{H}}_{\text{even}}\cup\vec{H}_{\text{odd}}$ is comprised of the original odd pseudopotentials $\vec{H}_{\text{odd}}$ and the free parameters $\tilde{\vec{H}}_{\text{even}}$. By construction, $\tilde{\vec{V}}$ yields the same energy as ${\vec{V}}$.
To reduce statistical noise, we tune $\tilde{\vec{H}}_{\text{even}}$ so that $\tilde{V}_{k}$ decay quickly by minimizing the cost function
${\cal F}(\tilde{\vec{H}}_{\text{even}})=\sum_{k=1}^{2q}[\tilde{V}_{k}(\tilde{\vec{H}}_{\text{even}})]^{2}(2k+1)^{\alpha},$ (19)
with $\alpha\geq 2$ to penalize large $k$ harmonics over small $k$foo We take $\alpha=4$ for the data presented in this work, but we find the final result to be largely insensitive of $\alpha$; see Fig. 3. After determining $\tilde{\vec{H}}^{\text{opt}}_{\text{even}}$ that minimizes ${\cal F}(\tilde{\vec{H}}_{\text{even}})$, we compute the refined values of the harmonics $\tilde{V}_{k}(\tilde{\vec{H}}^{\text{opt}}_{\text{even}})$. Using them to compute the energy Eq. (7) yields the same value as using the original $V_{k}=\tilde{V}_{k}({\vec{H}}_{\text{even}})$ but with less noise. The optimization procedure can be done semi-analytically, e.g., with Mathematica, and is not numerically expensive for any system size.
We note that Eq. (7) assumes spherically symmetric states. Nevertheless, the refined harmonics $\tilde{\vec{V}}$ can be cast into a real-space form $\tilde{V}(\theta)$ with Eq. (6) that can be used to compute the energy of non-zero angular momentum states within Monte Carlo simulations.